Find the angular momentum if postion vector is perpendicular with 30Ns Momentum of |τ|=0.5m
1
Expert's answer
2019-12-09T08:54:38-0500

The angular momentum is "\bold L = \bold r \times \bold p", and its absolute value is "L = |\bold r| |\bold p| \sin \theta" , where "\theta" is the angle between "\bold r" and "\bold p". Given that "\theta = 90^{\circ}", "L = 0.5 m \cdot 30 N s \cdot 1 = 15 N \cdot m \cdot s".